OpenStudy (anonymous):

Virus growth within populations - 20% of people who have the virus this year will contract it next year. - 80% of people who have the virus this year will NOT contract it next year. - 40% of people who have NOT had the virus this year will contract it next year. -60% of people who have NOT had the virus this year will NOT contract it next year. Given in a certain population 20000 out of 100000, use the above model to investigate the growth or otherwise of the flu virus in the population.

OpenStudy (anonymous):

\[\left[ \left[\begin{matrix}0.2 & 0.4 \\ 0.8 & 0.6\end{matrix}\right] \right]\]

OpenStudy (anonymous):

Is that the matrix which represents that question? with Contract and No contract as the rows and Virus and No Virus as the columns